# Penrose’s eight-conic theorem via Penrose’s eight-quadric theorem

### v4

@@ 1 Introduction
It is fair to say that neither Pappus’s theorem nor its dual are obvious. A natural question is whether there is a general pattern to theorems like Pappus’s, that could shed light on Pappus’s theorem, making it more intuitive and easier to discover.
Blaise Pascal (1623–1662), at age 16, discovered a remarkable generalization of Pappus’s theorem, known as Pascal’s mystic hexagram [DC20]. The initial line pair of Pappus’s theorem, as well as the initial point pair of the dual Pappus theorem, is a degeneration of a conic section. After replacing the pair of lines by an arbitrary conic, the rest of the statement of the theorem remains true (Figure 2).
@@
In 1848, George Salmon published a textbook on conic sections, which includes a further generalization of the Pascal–Brianchon theorem [Sal17, §264], rediscovered by Evelyn, Money-Coutts, and Tyrrell in 1974 [EMCT74, §2.3]. The three point pairs $AA^{\prime}$, $BB^{\prime}$, $CC^{\prime}$ of Pascal’s theorem are degenerations of thin ellipses that are in double contact with the initial conic. And the lines $AB^{\prime}$ and $A^{\prime}B$ are the limit of a pair of common tangents between the two thin ellipses collapsing to $AA^{\prime}$ and $BB^{\prime}$. After replacing the three point pairs by three conics in double contact with a fourth, the remainder of Pappus’s theorem statement still holds (Figure 3).
@@
The article is structured as follows. Section 2 provides the necessary background information on conics and quadrics. In Section 3, we present the main theorems: the eight-conic theorem and its 3D counterpart, the eight-quadric theorem. The proof of the eight-quadric theorem is given in Section 3.4, with the eight-conic theorem derived as a corollary through an extrusion process explained in Section 4. Section 5 illustrates examples of the eight-quadric configuration, showing how its 2D slices or projections yield classical theorems in the projective plane. The final sections of the article provide references to related works and future perspectives.
@@ 2.2.2 Some contact lemmas
On the other hand, if the chords of contact $p,q,r$ all meet in a point, then the span of $p^{2},q^{2},r^{2}$ contains only line pairs through this common point, so we have the following lemma:
@@
The following extended notation will be used below in Sec. 3.3.
2.3 Notation for complete quadrics
Every quadric has a point-wise, or primary, aspect as well as a plane-wise, or dual, aspect. We write the primary (dual) form of a quadric as $\mathbf{{Q}}$ ($\mathbf{{Q}}^{*}$). When does the pair $(\mathbf{{Q}},\mathbf{{Q}}^{*})$ represent a quadric? When either is regular, then so is the other, and the relationship is given by $\mathbf{{Q}}\mathbf{{Q}}^{*}=\lambda Id$ for $\lambda\neq 0$. Otherwise, both are degenerate and satisfy the relationship $\mathbf{{Q}}\mathbf{{Q}}^{*}=\lambda Id$ for $\lambda=0$, which can be seen by taking limits of regular quadrics. Then a necessary condition for $(\mathbf{{Q}},\mathbf{{Q}}^{*})$ to represent a quadric is $\mathbf{{Q}}\mathbf{{Q}}^{*}=\lambda Id$. If in addition both are not zero, we call the pair a complete quadric. In discussions limited to regular quadrics, we usually refer to a quadric simply in its point-wise form as $\mathbf{{Q}}$, since $\mathbf{{Q}}^{*}$ is uniquely determined. When degenerate quadrics play a role, we employ the full notation introduced here.
All the above also applies to conics, if plane-wise is replaced by line-wise.
@@ 3.2 The face figure of a Penrose conic configuration
In Fig. 5 (left), consider two conics corresponding to opposite nodes in a face of the cube, for example, $\mathbf{{T}}^{1}$ and $\mathbf{{T}}^{3}$. Then by (6) they can be written as $\mathbf{{T}}^{1}=\mathbf{{S}}_{0}+pp^{T}$ and $\mathbf{{T}}^{3}=\mathbf{{S}}_{0}+qq^{T}$ with $p,q$ being the chords of contact between $\mathbf{{S}}_{0}$ and $\mathbf{{T}}^{i}$. Hence $\mathbf{{T}}^{1}-\mathbf{{T}}^{3}=pp^{T}-qq^{T}=\frac{1}{2}(de^{T}+ed^{T})$ with $d=p+q$ and $e=p-q$ being two lines in the pencil spanned by $\mathbf{{T}}^{1}$ and $\mathbf{{T}}^{3}$. Thus, the chords of contact meet at the carrying point of the line pair $ed$. By Lemma 2.1, there are at most 3 line pairs in a regular pencil of conics. Hence, the set of conics that are in double contact with $\mathbf{{T}}^{1}$ and $\mathbf{{T}}^{3}$ splits into at most 3 families such that condition (ii) of the 8-conic theorem holds automatically for all conics $\mathbf{{S}}_{0},\mathbf{{S}}_{2}$ in the same family. All these conics can be explicitly written in terms of $d,e$ and $\mathbf{{S}}_{0}$ (see [Sal17, p.262]).
@@ 3.3 Penrose’s 8-conic theorem
The extrusion needed for the 7-configurations encountered in the following proofs is provided by the following lemma, whose proof appears below in Sect. 4.
@@ 3.4.1 The proof
We also assume for the moment that $o$ is not incident to $O$. This is equivalent to assuming that the 7 quadrics of the configuration do not all share a common contact element. We postpone to Sect. 4.1 the proof that this condition is not necessary, since it depends on elements of the proof of Lemma 3.1.
@@ 3.4.3 Refinement of Penrose 8-quadric theorem
4 Proof of Extrusion Lemma
@@
4.1 Proof of the 8-quadric theorem with a common contact element
@@
5 Examples of 8-quadric configurations
@@
The above examples show the following common structure: The three quadrics $\mathbf{{T}}^{1},\mathbf{{T}}^{2},\mathbf{{T}}^{3}$ give rise to 6 degenerate quadrics either by pairwise intersections (Examples 5.1 and 5.3) or by pairwise envelopes (Example 5.2). Depending on the choice of such a triplet $\mathbf{{S}}_{1},\mathbf{{S}}_{2},\mathbf{{S}}_{3}$, one from each pair, the final quadric $\mathbf{{T}}^{0}$ is either regular or a pair of points (or planes). This is summarized in Figure 12.
@@
6 Related research and outlook for future work
@@
We recently became aware of an independent discovery of the 8-conic theorem in [Eid13], Sect. 9.2.
@@
The research leading to this article has also produced two companion articles which provide a different perspective on the material. [CGN] presents a proof of the 8-conic theorem working within the 5D parameters space of conics. [ACG+] does the same for the 8-quadric theorem within the 9D space of quadrics. The proofs have a similar structure to each other, but the advantages of the 3D parameter space approach over the 2D approach become clear in the final stages of the proofs.
This article and [ACG+] are the result of a collaboration with Roger Penrose that emerged after the authors of [CGN] contacted him to report on their results. A collaboration ensued that led, among other results, to the successful reconstruction of the original 3D quadric proof of Roger Penrose in the two forms mentioned above.
@@
The final author recalls that as an undergraduate he first recognized the cube structure of the theorem from examples in which most of the conic sections are degenerate. He then developed a method for gluing together several such cubes to generate new cubes that contain more regular conics. In this way, he finally arrived at the formulation of the general theorem. Furthermore, the possibility of gluing several cubes together is also interesting to obtain further theorems. For example, this method can be used to extend Poncelet’s theorem for triangles to general polygons. Another example is Neville’s theorem about three ellipses that have a common focal point in pairs: Here, a first application of the eight-conic theorem can be used to show that all three ellipses must have double contact with another conic. This then serves as a starting point for further cubes, attached to the first one. We are currently working on examples of this type of iterated application of the theorem.

### v5

@@ 1 Introduction
It is fair to say that neither Pappus’s theorem nor its dual are obvious. A natural question is whether there is a general pattern to theorems like Pappus’s that could shed light on Pappus’s theorem, making it more intuitive and easier to discover.
Blaise Pascal (1623–1662), at age 16, discovered a remarkable generalization of Pappus’s theorem, known as Pascal’s mystic hexagram [DC20]. The initial line pair of Pappus’s theorem, as well as the initial point pair of the dual Pappus theorem, is a degeneration of a conic section. After replacing the pair of lines with an arbitrary conic, the rest of the statement of the theorem remains true (Figure 2).
@@
In 1848, George Salmon published a textbook on conic sections, which includes a further generalization of the Pascal–Brianchon theorem [Sal17, §264], rediscovered by Evelyn, Money-Coutts, and Tyrrell in 1974 [EMCT74, §2.3]. The three point pairs $AA^{\prime}$, $BB^{\prime}$, $CC^{\prime}$ of Pascal’s theorem are degenerations of thin ellipses that are in double contact with the initial conic. And the lines $AB^{\prime}$ and $A^{\prime}B$ are the limit of a pair of common tangents between the two thin ellipses collapsing to $AA^{\prime}$ and $BB^{\prime}$. After replacing the three point pairs with three conics in double contact with a fourth, the remainder of Pappus’s theorem statement still holds (Figure 3).
@@
The article is structured as follows. Section 2 provides the necessary background information on conics and quadrics. In Section 3, we present the main theorems: the eight-conic theorem and its 3D counterpart, the eight-quadric theorem. The proof of the eight-quadric theorem is given in Section 3.4, with the eight-conic theorem derived as a corollary through an extrusion process explained in Section 3.4.4. Section 3.5 illustrates examples of the eight-quadric configuration, showing how its 2D slices or projections yield classical theorems in the projective plane. Section 4 presents parametrized formulas for the conics of the combinatorial cube in terms of a minimal set of initial data. Section 5 of the article provides a brief account of related research and of the genesis of this article. Section 6 considers possible directions for further research.
@@ 2.2.2 Some contact lemmas
On the other hand, if the chords of contact $p,q,r$ all meet in a point, then the span of $pp^{T},qq^{T},rr^{T}$ contains only line pairs through this common point, so we have the following lemma:
@@
2.3 Notation
Before we turn to the proof of the theorems, we introduce some notation used in the sequel. First, an alternative notation for degenerate conics, used in Sec. 3.2 and Sec. 4. Secondly, notation for so-called complete quadrics, required to represent non-regular conics and quadrics, and used in Sec. 3.3.
2.3.1 Abbreviated notation
Until now we have written a line pair formed by the two lines $p$ and $q$ directly as a matrix via the symmetrized outer project $pq^{T}+qp^{T}$, and the double line of $p$ as $pp^{T}$. $p$ and $q$ can also be interpreted as the zero sets of the linear polynomials
$p(x,y,z):=ax+by+cz,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ q(x,y,z):=a^{\prime}x+b^{\prime}y+c^{\prime}z$
Their product is the quadratic polynomial
$r(x,y,z):$ $=p(x,y,z)q(x,y,z)$
$=aa^{\prime}x^{2}+bb^{\prime}y^{2}+cc^{\prime}z^{2}+(ab^{\prime}+a^{\prime}b)xy+(ac^{\prime}+a^{\prime}c)xz+(bc^{\prime}+b^{\prime}c)yz$
$=(x,y,z)^{T}(pq^{T}+qp^{T})(x,y,z)$
The final equality shows that $r(x,y,z)=0$ represents the same line pair as $(pq^{T}+qp^{T})$. Setting $q=p$ yields an expression for the double line $p^{2}$ which agrees with the matrix $pp^{T}$. We introduce the expressions $\mathbf{pq}$ and $\mathbf{p}^{2}$ as the abbreviated notation for the line pair, resp., the double line. We also write $\mathbf{p}$ for the line itself (not a double line); although it is not a conic, it can be used to generate conics within the ring of polynomials in the variables $\{x,y,z\}$. We continue to use $p$ to denote a line in $\mathbb{R}P^{2}$ when the membership in the polynomial ring is not highlighted.
Let $A$ and $B$ be two points. Similar remarks lead to the abbreviated notation $\mathbf{AB}$, resp., $\mathbf{A}^{2}$ for the point pair formed by $A$ and $B$, resp., the double point $A$.
Finally, in the sequel we also use $\mathbf{{X}}$ to represent the conic or quadric with matrix $X$. This ensures that bold-face font signals a conic, whether as $\mathbf{{X}}$ or the abbreviated forms $\mathbf{pq}$ and $\mathbf{p}^{2}$.
The advantage of the abbreviated notation is that it gives access to the algebraic structure of the ring of polynomials when evaluating linear combinations of conics involving degenerate conics, see for example Sec. 3.2. It was introduced by Salmon [Sal17].
2.3.2 Notation for complete quadrics
Every quadric has a point-wise, or primary, aspect as well as a plane-wise, or dual, aspect. We write the primary (dual) form of a quadric as $\mathbf{{Q}}$ ($\mathbf{{Q}}^{*}$). When does the pair $(\mathbf{{Q}},\mathbf{{Q}}^{*})$ represent a quadric? When either is regular, then so is the other, and the relationship is given by $\mathbf{{Q}}\mathbf{{Q}}^{*}=\lambda Id$ for $\lambda\neq 0$. Otherwise, both are degenerate and satisfy the relationship $\mathbf{{Q}}\mathbf{{Q}}^{*}=\lambda Id$ for $\lambda=0$, which can be seen by taking limits of regular quadrics. Then a necessary condition for $(\mathbf{{Q}},\mathbf{{Q}}^{*})$ to represent a quadric is $\mathbf{{Q}}\mathbf{{Q}}^{*}=\lambda Id$. If in addition both are not zero, we call the pair a complete quadric. In discussions limited to regular quadrics, we usually refer to a quadric simply in its point-wise form as $\mathbf{{Q}}$, since $\mathbf{{Q}}^{*}$ is uniquely determined. When degenerate quadrics play a role, as in Sec. 3.3, we employ the full notation introduced here.
All of the above also applies to conics, if plane-wise is replaced by line-wise.
@@ 3.2 The face figure of a Penrose conic configuration
In Fig. 5 (left), consider two conics corresponding to opposite nodes in a face of the cube, for example, $\mathbf{{T}}^{1}$ and $\mathbf{{T}}^{3}$. Then by (6) they can be written as $\mathbf{{T}}^{1}=\mathbf{{S}}_{0}+\mathbf{q}^{2}$ and $\mathbf{{T}}^{3}=\mathbf{{S}}_{0}+\mathbf{q}^{2}$ with $\mathbf{p}$ and $\mathbf{q}$ being the chords of contact between $\mathbf{{S}}_{0}$ and $\mathbf{{T}}^{i}$. Using abbreviated notation:
$\mathbf{{T}}^{1}-\mathbf{{T}}^{3}=\mathbf{p}^{2}-\mathbf{q}^{2}=(\mathbf{p}+\mathbf{q})(\mathbf{p}-\mathbf{q})=:\mathbf{de}$
where $\mathbf{d}:=\mathbf{p}+\mathbf{q}$ and $\mathbf{e}=\mathbf{p}-\mathbf{q}$ is a line pair in the pencil spanned by $\mathbf{{T}}^{1}$ and $\mathbf{{T}}^{3}$. Hence $\mathbf{p},\mathbf{q},\mathbf{d}$ and $\mathbf{e}$ have a common point. Consult Fig. 5, right.
By Lemma 2.1, there are at most 3 line pairs in a regular pencil of conics. Hence, the set of conics that are in double contact with $\mathbf{{T}}^{1}$ and $\mathbf{{T}}^{3}$ splits into at most 3 families such that condition (ii) of the 8-conic theorem holds automatically for all conics $\mathbf{{S}}_{0},\mathbf{{S}}_{2}$ in the same family. All these conics can be explicitly written in terms of $\mathbf{p}$, $\mathbf{q}$, and $\mathbf{{S}}_{0}$ (see [Sal17, §287]).
@@ 3.3 Penrose’s 8-conic theorem
The extrusion needed for the 7-configurations encountered in the following proofs is provided by the following lemma, whose proof appears below in Sect. 3.4.4.
@@ 3.4.1 The proof
We also assume for the moment that $o$ is not incident to $O$. This is equivalent to assuming that the 7 quadrics of the configuration do not all share a common contact element. We postpone to Sect. 3.4.5 the proof that this condition is not necessary, since it depends on elements of the proof of Lemma 3.1.
@@ 3.4.3 Refinement of Penrose 8-quadric theorem
3.4.4 Proof of Extrusion Lemma
@@
3.4.5 Proof of the 8-quadric theorem with a common contact element
@@
3.5 Examples of 8-quadric configurations
@@
The above examples show the following common structure: The three quadrics $\mathbf{{T}}^{1},\mathbf{{T}}^{2},\mathbf{{T}}^{3}$ give rise to 6 degenerate quadrics either by pairwise intersections (Examples 3.1 and 3.3) or by pairwise envelopes (Example 3.2). Depending on the choice of such a triplet $\mathbf{{S}}_{1},\mathbf{{S}}_{2},\mathbf{{S}}_{3}$, one from each pair, the final quadric $\mathbf{{T}}^{0}$ is either regular or a pair of points (or planes). This is summarized in Figure 12.
@@
4 Algebraic proof of the Penrose theorems
We now turn to a purely algebraic proof of Penrose’s 8-conic theorem, which presents explicit parametrized formulas for all 8 conics of a Penrose cube, starting from $\mathbf{{S}}_{0}$ and its 3 chords of contact with $\mathbf{{T}}^{1},\mathbf{{T}}^{2}$, and $\mathbf{{T}}^{3}$, respectively. It presents these formulas without detailed derivations. It focuses on the case of point-wise conics; similar formulas exist also for line-wise conics.
4.1 The proof
The proof makes repeated use of the fact that two point-wise conics are in double contact with each other if and only if their pencil contains a double line. That is, there exists $a$ and $b$ such that
$a\,\mathbf{{C}}_{1}+b\,\mathbf{{C}}_{2}=\mathbf{m}^{2}$ (45)
where ${m}$ is the joining line of the contact points. When we have established this relation for a pair of conics, we have proved that the two conics are in double contact.
We start with a conic $\mathbf{{S}}_{0}$ and a line $\mathbf{p}$. Consider the sum:
$\mathbf{{T}}^{1}=\mathbf{{S}}_{0}+\mathbf{p}^{2}$ (46)
Then $\mathbf{{T}}^{1}$ and $\mathbf{{S}}_{0}$ are in double contact because we have $\mathbf{{T}}^{1}-\mathbf{{S}}_{0}=\mathbf{p}^{2}$, the square of a line as stated above.
In the same way, we form a conic $\mathbf{{T}}^{2}$ and $\mathbf{{T}}^{3}$ from $\mathbf{{S}}_{0}$ and lines $\mathbf{q}$ and $\mathbf{r}$:
$\mathbf{{T}}^{2}$ $=$ $\mathbf{{S}}_{0}+\mathbf{q}^{2}$
$\mathbf{{T}}^{3}$ $=$ $\mathbf{{S}}_{0}+\mathbf{r}^{2}$
There is now the important property that the conic $\mathbf{{S}}_{0}$ can move along two conics in double contact with it, while staying in double contact with both. The family of such conics can be expressed using $\mathbf{{S}}_{0}$, the chords of contact $\mathbf{q}$ and $\mathbf{r}$, and a free parameter $a$:
$\mathbf{{S}}_{1}$ $=$ $(1-a^{2})\,\mathbf{{S}}_{0}+\mathbf{q}^{2}+2\,a\,\mathbf{qr}+\mathbf{r}^{2}$ (47)
And we find the conics in double contact with $\mathbf{{T}}^{1}-\mathbf{{T}}^{3}$ and $\mathbf{{T}}^{1}-\mathbf{{T}}^{2}$
$\mathbf{{S}}_{2}$ $=$ $(1-b^{2})\,\mathbf{{S}}_{0}+\mathbf{p}^{2}+2\,b\,\mathbf{pr}+\mathbf{r}^{2}$
$\mathbf{{S}}_{3}$ $=$ $(1-c^{2})\,\mathbf{{S}}_{0}+\mathbf{p}^{2}+2\,c\,\mathbf{pq}+\mathbf{q}^{2}$
where $b$ and $c$ are free constants.
That $\mathbf{{S}}_{3}$ is in double contact with $\mathbf{{T}}^{1}$ we see by the fact that a linear combination of them gives a double line:
$\mathbf{{S}}_{3}-(1-c^{2})\mathbf{{T}}^{1}=(c\mathbf{p}+\mathbf{q})^{2}$ (48)
where $c\mathbf{p}+\mathbf{q}$ is the chord of contact. The symmetry shows that there are such contacts between the other conics as well.
The last to show is that the three conics $\mathbf{{S}}_{1}$, $\mathbf{{S}}_{2}$ and $\mathbf{{S}}_{3}$ are in double contact with an eighth conic. This one is totally determined by the other elements in the structure, and the formula is given by:
$\mathbf{{T}}^{0}$ $=$ $(1-a^{2}-b^{2}-2abc-c^{2})\,\mathbf{{S}}_{0}$
$+$ $(1-a^{2})\,\mathbf{p}^{2}+(1-b^{2})\,\mathbf{q}^{2}+(1-c^{2})\,\mathbf{r}^{2}$
$+$ $2(a+b\,c)\,\mathbf{qr}+2(b+ac)\,\mathbf{pr}+2(c+ab)\,\mathbf{pq}$
We find that $\mathbf{{T}}^{0}$ and $\mathbf{{S}}_{1}$ are in double contact because we have:
$\begin{split}&(1-a^{2})\,\mathbf{{T}}^{0}-(1-a^{2}-b^{2}-2abc-c^{2})\,\mathbf{{S}}_{1}\\ &=\{(1-a^{2})\,\mathbf{p}+(c+ab)\,\mathbf{q}+(b+ac)\,\mathbf{r}\}^{2}\end{split}$ (50)
where
$(1-a^{2})\,\mathbf{p}+(c+ab)\,\mathbf{q}+(b+ac)\,\mathbf{r}$
is the line through the contact points of $\mathbf{{T}}^{0}$ and $\mathbf{{S}}_{1}$. From the symmetry, we see that $\mathbf{{S}}_{2}$ and $\mathbf{{S}}_{3}$ are also in double contact with $\mathbf{{T}}^{0}$.
This proof also applies to quadrics. One has only to replace double contact with ring contant, and the line pairs and double lines with plane pairs, resp., double planes.
5 Related research and a brief history
@@
We recently became aware of an independent discovery of the 8-conic theorem [Eid13], Sect. 9.2. (English translation here.) The author Morten Eide has in the meantime become a co-author of this article and contributed Sec. 4.
The impetus for this article came from a podcast featuring Roger Penrose [Har20], in which he describes formulating the 8-conic theorem in the early 1950’s, but not having found the time to publish it.
In 2024 the authors of [CGN] contacted Roger Penrose to report their proof of the theorem. A collaboration ensued, in which Roger Penrose shared details of his original approaches to the proof of the theorem. He recalled that as an undergraduate he first recognized the cube structure of the theorem from examples in which most of the conic sections are degenerate. He then developed a method for gluing together several such cubes to generate new cubes that contain more regular conics. In this way, he finally arrived at the formulation of the general theorem. Later, as a graduate student, he discovered an approach to prove the 8-conic theorem using an analogous 8-quadric theorem in 3-D. It is this approach that is reconstructed in the proof in Sec. 3.
@@
Related research has produced two companion articles that provide a different perspective on the material. [CGN] presents a proof of the 8-conic theorem working within the 5-D parameters space of conics. [ACG+] does the same for the 8-quadric theorem within the 9-D space of quadrics. The proofs have a similar structure to each other; the advantages of the 3-D parameter space approach over the 2-D approach become clear in the final stages of the proofs.
6 Outlook for future work
@@
The possibility of gluing several Penrose cubes together, described above, remains an interesting option to obtain further theorems. For example, this method can be used to extend Poncelet’s theorem for triangles to general polygons. Another example is Neville’s theorem about three ellipses that have a common focal point in pairs: Here, a first application of the eight-conic theorem can be used to show that all three ellipses must have double contact with another conic. This then serves as a starting point for further cubes, attached to the first one. We are currently working on examples of this type of iterated application of the theorem.